A certain weak base has a Kb of 8.40 × 10-7. What concentration of this base will produce a pH of 10.09?
Answer #1

As pH 10.09, and using the formula, pH + pOH = 14

pOH = 14 - 10.09 = 3.91

Using the formula for pH of a weak base

Kb = 8.4× 10-7

pKb = - log(8.4×10-7) = 6.07

pOH = 1/2(pKb - logC) = 1/2(6.07- logC) = 3.91

LogC = 6.07 - 2×3.91 = -1. 75

So C = concentration= 10-1.75 = 0.018M